OpenStudy (anonymous):

A cyclist rides her bike at a rate of 8 meters per second. What is this rate in kilometers per hour? How many kilometers will the cyclist travel in 5 hours? Do not round your answers.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

8\[\Large 8 \frac{ m }{ s }=8*\frac{ 1000m (=1km) }{ 3600s (=1h) }\]

OpenStudy (anonymous):

8 km/h

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is that the answer for the first question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh actually it's the other way around. \[8*\frac{ 3600 }{ 1000 }\]

OpenStudy (anonymous):

125?